Marin Company’s net income for 2020 is $54,900. The only potentially dilutive securities outstanding were 800 options issued during 2019, each exercisable for one share at $6. None has been exercised, and 10,200 shares of common were outstanding during 2020. The average market price of Marin’s stock during 2020 was $25.

(b) Assume the same facts as those assumed for part (a), except that the 800 options were issued on October 1, 2020 (rather than in 2019). The average market price during the last 3 months of 2020 was $25. (Round answer to 2 decimal places, e.g. $2.55.)

Diluted earnings per share

$
Answer #1

Diluted options = 800-(800*6/25) = 608*3/12 = 152

Diluted earning per share = Adjusted net income/Adjusted shares = 54900/10352 = 5.30 per share